# Progressive Contrastive Learning with Multi-Prototype for Unsupervised Visible-Infrared Person Re-identification

### v1

@@ Abstract
Unsupervised visible-infrared person re-identification (USVI-ReID) aims to match specified people in infrared images to visible images without annotation, and vice versa. USVI-ReID is a challenging yet under-explored task. Most existing methods address the USVI-ReID problem using cluster-based contrastive learning, which simply employs the cluster center as a representation of a person. However, the cluster center primarily focuses on shared information, overlooking disparity. To address the problem, we propose a Progressive Contrastive Learning with Multi-Prototype (PCLMP) method for USVI-ReID. In brief, we first generate the hard prototype by selecting the sample with the maximum distance from the cluster center. This hard prototype is used in the contrastive loss to emphasize disparity. Additionally, instead of rigidly aligning query images to a specific prototype, we generate the dynamic prototype by randomly picking samples within a cluster. This dynamic prototype is used to retain the natural variety of features while reducing instability in the simultaneous learning of both common and disparate information. Finally, we introduce a progressive learning strategy to gradually shift the model’s attention towards hard samples, avoiding cluster deterioration. Extensive experiments conducted on the publicly available SYSU-MM01 and RegDB datasets validate the effectiveness of the proposed method. PCLMP outperforms the existing state-of-the-art method with an average mAP improvement of 3.9%. The source codes will be released.
1 Introduction
Figure 1: Illustration for the core idea of our proposal. Left: Initial sample distribution and movement trends during optimization. Right: Sample distribution upon completion of optimization. CPCL denotes Centroid Prototype Contrastive Learning, HPCL for Hard Prototype Contrastive Learning, and DPCL for Dynamic Prototype Contrastive Learning.
@@ 1 Introduction
Visible-infrared person re-identification (VI-ReID) aims at matching the same person captured in one modality with their counterparts in another modality Ye et al. (2021); Gong et al. (2024); Yang et al. (2024). With the development of deep learning, VI-ReID has achieved remarkable advancements Yu et al. (2023); Kim et al. (2023); Feng et al. (2023); Fang et al. (2023). However, the development of existing VI-ReID methods is still limited due to the requirement for expensive-annotated training data Pang et al. (2022). To mitigate the problem of annotating large-scale cross-modality data, some semi-supervised VI-ReID methods Wang et al. (2022); Shi et al. (2023); Yang et al. (2023b) are proposed to learn modality-invariant and identity-related discriminative representations by utilizing both labeled and unlabeled data. For this purpose, OTLA Wang et al. (2022) proposed an optimal-transport label assignment mechanism to assign pseudo-labels for unlabeled infrared images while ignoring how to calibrate noise pseudo-labels. DPIS Shi et al. (2023) integrates two pseudo-labels generated by distinct models into a hybrid pseudo-label for unlabeled infrared data, but it makes the training process more complex. Although these methods have gained promising performances, they still rely on a certain number of manual-labeled data.
@@
USVI-ReID is gaining prominence and several methods Yang et al. (2022a); Zhang et al. (2023b); Yang et al. (2023c); Wu and Ye (2023) have proposed to tackle the issues of expensive visible-infrared annotation through contrastive learning. These methods create two modality-specific memories, one for visible features and the other for infrared features. During training, these methods consider the memory center as a centroid prototype and minimize the contrastive loss across the features of query images and centroid prototype. Then, these methods aggregate the corresponding centroid prototypes based on similarity. However, the centroid prototype only stores the common information of each person, neglecting the diverse information. Just like a normal distribution, to better reflect the data distribution of a dataset, we need not only the mean but also the variance. Additionally, contrastive loss based on the centroid prototype may blur the boundaries across different clusters as illustrated in the top row of Fig. 1.
@@
In this paper, we rethink the design of the prototype used in contrastive learning for USVI-ReID. Firstly, we design a Hard Prototype Contrastive Learning (HPCL) to mine diverse yet meaningful information and improve the boundaries across different clusters. In contrast to traditional contrastive learning methods, we choose the hard samples to serve as the hard prototype as shown in the middle row of Fig. 1. In other words, the hard prototype is the one that is farthest from the memory center. The hard prototype encompasses distinctive information. Furthermore, we introduce the concept of Dynamic Prototype Contrastive Learning (DPCL), we randomly select samples from each cluster to serve as the dynamic prototype. DPCL effectively accounts for the intrinsic diversity within clusters, enhancing the model’s adaptability to varying data distributions. The dynamic prototype offers more comprehensive representations and relationships compared with memory samples overlapping, as shown in the bottom low of Fig. 1. Early clustering results are unreliable, and utilizing hard samples at this stage may lead to cluster degradation. Therefore, we introduce progressive contrastive learning to gradually focus on hard samples.
@@
  • We propose a progressive contrastive learning with multi-prototype method for the USVI-ReID. We reconsider the design of prototypes in progressive contrastive learning to ensure that the model stably captures both commonality and diversitiy.
  • We propose Hard Prototype Contrastive Learning for mining distinctive yet significant information, and Dynamic Prototype Contrastive Learning for preserving the intrinsic diversity in sample features.
  • The experimental results on the SYSU-MM01 and RegDB datasets with superior performance against the state-of-the-art methods demonstrate the effectiveness of the proposed method.
@@ 2.1 Supervised Visible-Infrared Person ReID
Visible-Infrared Person ReID (VI-ReID) has drawn much attention in recent years Ye et al. (2021); Yang et al. (2022b). Many VI-ReID works focused on mitigating huge semantic gaps across modalities have made advanced progress, which can be classified into two primary classes based on their different aligning ways: image-level alignment and feature-level alignment. The image-level alignment methods focus on reducing cross-modality gaps by modality translation. Some GAN-based methods Wang et al. (2020, 2019) are proposed to perform style transformation for aligning cross-modality images. However, the generated images unavoidably contain noise. Therefore, X-modality Li et al. (2020) and its promotion Zhang et al. (2021); Wei et al. (2021) align cross-modality images by introducing a middle modality. Mainstream feature-level alignment methods Liu et al. (2022); Zhang et al. (2022); Ye et al. (2020); Sun et al. (2022); Wu et al. (2021) focus on minimizing cross-modality gaps by finding a modality-shared feature space. To this end, SAAI Fang et al. (2023) proposes a Semantic Alignment and Affinity Inference framework, which aligns the latent semantic part features by shared prototypes, enabling the aggregation of latent semantic part features with the same semantic information across both modalities. However, the advanced performances of the above methods build on large-scale human-labeled cross-modality data, which are quite time-consuming and expensive, thus hindering the fast application of these methods in real-scenes.
Figure 2: Framework of our PCLMP. The framework consists of two stages: the first stage employs contrastive learning with centroid prototypes to learn well-discriminative representation, and the second stage introduces contrastive learning with hard and dynamic prototypes to further focus on hard samples.
@@ 2.2 Unsupervised Single-Modality Person ReID
The existing unsupervised single-modality person ReID methods can be roughly divided into two classes: Unsupervised domain adaption methods (UDA), which try to leverage the knowledge transferred from labeled source domain to improve performance Ge et al. (2020a, b), and Purely unsupervised methods (USL), which directly train a USL-ReID model on the unlabeled target domain Zhang et al. (2023b); Cho et al. (2022). Compared with the UDA methods, the USL methods are more challenging. Recently, cluster-contrast learning Dai et al. (2022) has achieved impressive performance by performing contrastive learning at the cluster level. However, cluster-contrast with a uni-proxy can be biased and confusing, which fails to accurately describe the information of a cluster. To this end, methods Zou et al. (2023); Wu et al. (2022) proposed maintaining multi-proxies for a cluster to adaptively capture different information within the cluster. The above methods are mainly proposed to solve the single-modality ReID task, but they are limited to solving the USL-VI-ReID task due to large cross-modality gaps.
@@ 2.3 Unsupervised Visible-Infrared Person ReID
Recently, unsupervised visible-infrared person ReID has attracted much attention due to the advantage of not relying on any data annotation. Some UDA methods Liang et al. (2021); Wang et al. (2022) use a well-annotated labeled source domain for pre-training to solve the USVI-ReID task. Some purely unsupervised methods Yang et al. (2022a); Wu and Ye (2023); Yang et al. (2023c, a) adopt contrastive learning to boost performance inspired by Dai et al. (2022), which mainly follow a two-step loop paradigm: generating pseudo-labels using the DBSCAN algorithm Ester et al. (1996) to create memory banks with clustering centers and establishing cross-modality correspondences based on these memory banks. However, pseudo-labels are often inaccurate and rigid, CCLNet Chen et al. (2023) leverages the text information from CLIP to afford greater semantic monitoring insights to compensate for the rigidity of pseudo-labels. Moreover, reliable cross-modality correspondences are vital to USVI-ReID, thus PGM Wu and Ye (2023) proposes a progressive graph matching framework to establish more reliable cross-modality correspondences. However, cluster centers mainly present common information while lacking distinctive information, which results in ambiguous cross-modality correspondences when meeting hard samplesYang et al. (2021, 2023d)
@@ 3 Method
3.1 Overview and Problem Formulation
We propose a Progressive Contrastive Learning with Multi-Prototype (PCLMP) method for USVI-ReID, which mainly contains an online backbone, momentum backbone, and progressive contrastive learning strategy with a centroid prototype, a hard prototype, and a dynamic prototype, as shown in Fig. 2. The online backbone is a standard network, updated through back-propagation. The momentum backbone mirrors the structure of the online backbone, updated through the weights of the online backbone. The clustering is used to generate pseudo labels for creating cluster-aware memory, and we employ DBSCAN Ester et al. (1996) for this clustering. We use progressive contrastive learning with multi-prototype to optimize the online backbone. PCLMP primarily focuses on representation learning, and we use PGM Wu and Ye (2023) to aggregate cross-modality memory.
Given a USVI-ReID dataset $D=\{V,R\}$, where $V=\{V_{i}\}^{N_{v}}_{i=1}$ represents the visible images and $R=\{R_{j}\}^{N_{r}}_{j=1}$ denotes the infrared images. $V_{i}$ and $R_{j}$ represent the set of images corresponding to the $i$-th and $j$-th class. $N_{v}$ and $N_{r}$ denote the number of visible and infrared clusters, respectively. $q^{v}$ and $q^{r}$ are visible and infrared query instance features. In the USVI-ReID task, the purpose is to train a deep neural network to obtain modality-invariant and identity-related features from images without any annotation.
@@ 3.2 Centroid Prototype Contrastive Learning
Following the USVI-ReID methods Yang et al. (2022a, 2023c); Shi et al. (2024), our method uses centroid prototype contrastive learning to optimize the online backbone $\phi_{0}$ in the first state, which includes memory initialization and optimization.
@@
$m^{v}_{i}=\frac{1}{|V_{i}|}\sum_{v\in V_{i}}\phi_{0}(v),$ (1)
$m^{r}_{j}=\frac{1}{|R_{j}|}\sum_{r\in R_{j}}\phi_{0}(r),$ (2)
where ||$\cdot$|| ⋅ | denotes the number of instances.
Optimization. During training, we update the two modality-specific memories by a momentum updating strategy Dai et al. (2022). We treat the memory center as a centroid prototype and optimize the feature extractor $\phi_{0}$ using contrastive learning with the centroid prototype, computed as:
$L^{v}_{CPCL}=-\log\frac{\exp\left(q^{v}\cdot{m_{+}^{v}}/\tau\right)}{\sum_{i=1}^{N_{v}}\exp\left(q^{v}\cdot{m_{i}^{v}}/\tau\right)},$ (3)
$L^{r}_{CPCL}=-\log\frac{\exp\left(q^{r}\cdot{m_{+}^{r}}/\tau\right)}{\sum_{j=1}^{N_{r}}\exp\left(q^{r}\cdot{m_{j}^{r}}/\tau\right)},$ (4)
$L_{CPCL}=L^{v}_{CPCL}+L^{r}_{CPCL},$ (5)
where $m_{+}^{v(r)}$ is the positive centroid prototype, denoting a query and the prototype shares the same pseudo label. The $\tau$ is a temperature hyper-parameter.
@@ 3.3 Hard Prototype Contrastive Learning
To ensure that the prototype effectively captures both the commonality and variation in a person, we devise a novel prototype for contrastive learning, which is referred to as Hard Prototype Contrastive Learning. HPCL is designed to provide a comprehensive understanding of personal characteristics, which benefits its handling of hard samples.
We use the online backbone $\phi_{0}$ to extract feature representations, and select $k$ samples that are farthest from the memory center as the hard prototype:
$hm^{v}_{i}=\underset{{{\forall v\in V_{i}}}}{\arg\max}\left\|\phi_{0}(v)-m^{v}_{i}\right\|,$ (6)
$hm^{r}_{j}=\underset{{{\forall r\in R_{j}}}}{\arg\max}\left\|\phi_{0}(r)-m^{r}_{j}\right\|.$ (7)
@@
$L_{HPCL}^{v}=-\log\frac{\exp\left(q^{v}\cdot{hm_{+}^{v}}/\tau\right)}{\sum_{i=1}^{N_{v}}\exp\left(q^{v}\cdot{hm_{i}^{v}}/\tau\right)},$ (8)
$L_{HPCL}^{r}=-\log\frac{\exp\left(q^{r}\cdot{hm_{+}^{r}}/\tau\right)}{\sum_{j=1}^{N_{r}}\exp\left(q^{r}\cdot{hm_{j}^{r}}/\tau\right)},$ (9)
$L_{HPCL}=L_{HPCL}^{v}+L_{HPCL}^{r},$ (10)
@@
$hm^{v}_{+(t)}=\alpha{hm}^{v}_{+(t-1)}+(1-\alpha)q^{v}_{t},$ (11)
$hm^{r}_{+(t)}=\alpha{hm}^{r}_{+(t-1)}+(1-\alpha)q^{r}_{t},$ (12)

### v2

# Learning Commonality, Divergence and Variety for Unsupervised Visible-Infrared Person Re-identification

@@ Abstract
Unsupervised visible-infrared person re-identification (USVI-ReID) aims to match specified people in infrared images to visible images without annotation, and vice versa. USVI-ReID is a challenging yet under-explored task. Most existing methods address the USVI-ReID problem using cluster-based contrastive learning, which simply employs the cluster center as a representation of a person. However, the cluster center primarily focuses on commonality, overlooking divergencies. To address the problem, we propose a Progressive Contrastive Learning with Hard and Dynamic Prototypes method for USVI-ReID. In brief, we first generate the hard prototype by selecting the sample with the maximum distance from the cluster center. We theoretically show that the hard prototype is used in the contrastive loss to emphasize divergence. Additionally, instead of rigidly aligning query images to a specific prototype, we generate the dynamic prototype by randomly picking samples within a cluster. The dynamic prototype is used to encourage the variety of features. Finally, we introduce a progressive learning strategy to gradually shift the model’s attention towards divergence and variety, avoiding cluster deterioration. Extensive experiments conducted on the publicly available SYSU-MM01 and RegDB datasets validate the effectiveness of the proposed method. Our method outperforms the existing state-of-the-art method with an average mAP improvement of 3.9%. The source codes will be released.
1 Introdction
@@
Visible-infrared person re-identification (VI-ReID) aims at matching the same person captured in one modality with their counterparts in another modality [1, 2, 3, 4]. It has recently gained attention in computer vision applications like video surveillance, security, and person analysis. With the development of deep learning, VI-ReID has achieved remarkable advancements [5, 6, 7, 8, 9]. However, the development of existing VI-ReID methods is still limited due to the requirement for expensive-annotated training data [10]. To mitigate the problem of annotating large-scale cross-modality data, some semi-supervised VI-ReID methods [11, 12, 13] are proposed to learn modality-invariant and identity-related discriminative representations by utilizing both labeled and unlabeled data. For this purpose, OTLA [11] proposed an optimal transport label assignment mechanism to assign pseudo-labels for unlabeled infrared images while ignoring how to calibrate noise pseudo-labels. DPIS [12] integrates two pseudo-labels generated by distinct models into a hybrid pseudo-label for unlabeled infrared data, but it makes the training process more complex. Although these methods have gained promising performances, they still rely on a certain number of manual-labeled data.
@@
Several USVI-ReID methods [14, 15, 16, 17] have proposed to tackle the issues of expensive visible-infrared annotation through contrastive learning. These methods create two modality-specific memories, one for visible features and the other for infrared features. During training, these methods consider the memory center as a prototype and minimize the contrastive loss across the features of query images and prototype. Then, these methods aggregate the corresponding prototypes based on similarity. However, the centroid prototype only stores the commonality of each person, neglecting the divergence, which causes the pseudo-labels generated by the cluster to be unreliable. Just like a normal distribution, to better reflect the data distribution of a dataset, we need not only the mean but also the variance. To this end, we pay attention to the cluster divergence, constraining the distance between the furthest points and the centroid to ensure cluster compactness.
@@
In this paper, we argue that an important aspect of contrastive learning for USVI-ReID, i.e. the design of the prototype, has so far been neglected, and propose progressive contrastive learning with hard and dynamic prototype (PCLHD) method for the USVI-ReID. Firstly, we design a Hard Prototype Contrastive Learning (HPCL) to mine divergent yet meaningful information. In contrast to traditional contrastive learning methods, we choose the hard samples to serve as the hard prototype. In other words, the hard prototype is the one that is farthest from the memory center. The hard prototype encompasses distinctive information. Furthermore, we introduce the concept of Dynamic Prototype Contrastive Learning (DPCL), we randomly select samples from each cluster to serve as the dynamic prototype. DPCL effectively accounts for the intrinsic variety within clusters, enhancing the model’s adaptability to varying data distributions. Early clustering results are unreliable, and utilizing hard samples at this stage may lead to cluster degradation. Therefore, we introduce progressive contrastive learning to gradually focus on hard samples.
@@
  • We propose a progressive contrastive learning with hard and dynamic prototype method for the USVI-ReID. We reconsider the design of prototypes in contrastive learning to ensure that the model stably captures commonality, divergence, and variety.
  • We propose Hard Prototype Contrastive Learning for mining divergent yet significant information, and Dynamic Prototype Contrastive Learning for preserving the intrinsic variety in sample features.
  • Experiments on SYSU-MM01 and RegDB datasets demonstrate the superiority of our method compared with existing USVI-ReID methods, and PCLHD generates higher-quality pseudo-labels than other methods.
@@ 2.1 Supervised Visible-Infrared Person ReID
Visible-Infrared Person ReID (VI-ReID) has drawn much attention in recent years [1, 18, 19, 20, 21, 22]. Many VI-ReID works focused on mitigating huge semantic gaps across modalities have made advanced progress, which can be classified into two primary classes based on their different aligning ways: image-level alignment and feature-level alignment. The image-level alignment methods focus on reducing cross-modality gaps by modality translation. Some GAN-based methods [23, 24] are proposed to perform style transformation for aligning cross-modality images. However, the generated images unavoidably contain noise. Therefore, X-modality [25] and its promotion [26, 27] align cross-modality images by introducing a middle modality. Mainstream feature-level alignment methods [28, 29, 30, 31, 32] focus on minimizing cross-modality gaps by finding a modality-shared feature space. However, the advanced performances of the above methods build on large-scale human-labeled cross-modality data, which are quite time-consuming and expensive, thus hindering the fast application of these methods in real-scenes.
Figure 1: Framework of our PCLHD. The framework consists of two stages: the first stage employs contrastive learning with centroid prototypes to learn well-discriminative representation, and the second stage introduces contrastive learning with hard and dynamic prototypes to further focus on divergence and variety.
@@ 2.2 Unsupervised Single-Modality Person ReID
The existing unsupervised single-modality person ReID methods can be roughly divided into two classes: Unsupervised domain adaption methods (UDA), which try to leverage the knowledge transferred from labeled source domain to improve performance [33, 34, 35, 36, 37], and fully unsupervised methods (USL), which directly train a USL-ReID model on the unlabeled target domain [15, 38]. Compared with the UDA methods, the USL methods are more challenging. Recently, cluster-contrast learning [39] has achieved impressive performance by performing contrastive learning at the cluster level. However, cluster-contrast with a uni-proxy can be biased and confusing, which fails to accurately describe the information of a cluster. To this end, methods [40, 41] proposed maintaining multi-proxies for a cluster to adaptively capture different information within the cluster. The above methods are mainly proposed to solve the single-modality ReID task, but they are limited to solving the USL-VI-ReID task due to large cross-modality gaps.
@@ 2.3 Unsupervised Visible-Infrared Person ReID
Unsupervised visible-infrared person ReID has attracted much attention due to the advantage of not relying on any data annotation. Some UDA methods [42, 11] use a well-annotated labeled source domain for pre-training to solve the USVI-ReID task. Some fully unsupervised methods [14, 17, 16, 43] adopt contrastive learning to boost performance inspired by [39], which mainly follow a two-step loop paradigm: generating pseudo-labels using the DBSCAN algorithm [44] to create memory banks with clustering centers and establishing cross-modality correspondences based on these memory banks. However, pseudo-labels are often inaccurate and rigid, CCLNet [45] leverages the text information from CLIP to afford greater semantic monitoring insights to compensate for the rigidity of pseudo-labels. Moreover, reliable cross-modality correspondences are vital to USVI-ReID, thus PGM [17] proposes a progressive graph matching framework to establish more reliable cross-modality correspondences. However, cluster centers mainly present common information while lacking distinctive information, which results in ambiguous cross-modality correspondences when meeting hard samples[46, 47].
@@ 3 Method
3.1 Problem Formulation and Overview
Given a USVI-ReID dataset $D=\{V,R\}$, where $V=\{V_{i}\}^{N_{v}}_{i=1}$ represents the visible images and $R=\{R_{j}\}^{N_{r}}_{j=1}$ denotes the infrared images. $V_{i}$ and $R_{j}$ represent the set of images corresponding to the $i$-th and $j$-th class. $N_{v}$ and $N_{r}$ denote the number of visible and infrared clusters, respectively. In the USVI-ReID task, the purpose is to train a deep neural network to obtain modality-invariant and identity-related features for matching pedestrian images with the same identity.
We propose a Progressive Contrastive Learning with Hard and Dynamic Prototype (PCLHD) method for USVI-ReID, which mainly contains an online encoder, momentum encoder, and progressive contrastive learning strategy with a centroid prototype, a hard prototype, and a dynamic prototype, as shown in Fig. 1. The online encoder is a standard network, updated through back-propagation. The momentum encoder mirrors the structure of the online encoder, updated through the weights of the online encoder. The clustering is used to generate pseudo labels for creating cluster-aware memory, and we employ DBSCAN [44] for clustering. PCLHD primarily focuses on representation learning, and we use PGM [17] to aggregate cross-modality memory.
@@ 3.2 Centroid Prototype Contrastive Learning
Following the USVI-ReID methods [14, 16, 48], we use centroid prototype contrastive learning to optimize the online encoder in the first state, which includes memory initialization and optimization.
@@
${cm}^{v}_{i}=\frac{1}{|V_{i}|}\sum_{v\in V_{i}}\phi_{0}(v),$ (1)
${cm}^{r}_{j}=\frac{1}{|R_{j}|}\sum_{r\in R_{j}}\phi_{0}(r),$ (2)
where ||$\cdot$|| ⋅ | denotes the number of instances belonging to specific cluster.
Optimization. During training, we update the two modality-specific memories by a momentum updating strategy [39]. We treat the memory center as a centroid prototype and optimize the feature extractor $\phi_{0}$ using contrastive learning with the centroid prototype, computed as:
$\mathcal{L}_{CPCL}^{v}=\frac{1}{N_{v}}\sum_{i\in N_{v}}\frac{-1}{|V_{i}|}\sum_{v\in V_{i}}\log{\frac{\text{exp}\left({\phi}_{0}(v)\cdot{cm}_{+}^{v}/\tau\right)}{\sum\limits_{j\in N_{v}}\text{exp}\left({{\phi}_{0}(v)}\cdot{cm}^{v}_{j}/\tau\right)}},$ (3)
$\mathcal{L}_{CPCL}^{r}=\frac{1}{N_{r}}\sum_{i\in N_{r}}\frac{-1}{|R_{i}|}\sum_{r\in R_{i}}\log{\frac{\text{exp}\left({\phi}_{0}(r)\cdot{cm}_{+}^{r}/\tau\right)}{\sum\limits_{j\in N_{r}}\text{exp}\left({{\phi}_{0}(r)}\cdot{cm}^{r}_{j}/\tau\right)}},$ (4)
$\mathcal{L}_{CPCL}=\mathcal{L}^{v}_{CPCL}+\mathcal{L}^{r}_{CPCL},$ (5)
where $cm_{+}^{v(r)}$ is the positive centroid prototype, denoting a query and the prototype shares the same identity. The $\tau$ is a temperature hyper-parameter.
@@ 3.3 Hard Prototype Contrastive Learning
To ensure that the prototype effectively captures divergence within a identity, we devise a novel hard prototype for contrastive learning, which is referred to as Hard Prototype Contrastive Learning. HPCL is designed to provide a comprehensive understanding of personal characteristics, which benefits its handling of hard samples. We use the online encoder $\phi_{0}$ to extract feature representations, and select $k$ samples that are farthest from the memory center as the hard prototype:
$hm^{v}_{i}=\underset{{{\forall v\in V_{i}}}}{\arg\max}\left\|\phi_{0}(v)-{cm}^{v}_{i}\right\|,$ (6)
$hm^{r}_{j}=\underset{{{\forall r\in R_{j}}}}{\arg\max}\left\|\phi_{0}(r)-{cm}^{r}_{j}\right\|.$ (7)
Theorem The information entropy of hard sample prototypes is greater than the information entropy of centroid prototypes, thereby preserving greater divergence within the hard memory.
Proof Given a set of features $\{\mathbf{z}_{1},\mathbf{z}_{2},\ldots,\mathbf{z}_{N_{c}}\}$ for class $c$. The entropy $H(\mathbf{cm}_{c})$ can be approximated by the entropy of the distribution of the sample means. Considering that $\mathbf{cm}_{c}$ is a convex combination of the sample features $\mathbf{z}_{i}$, we have:
$H\left(\mathbf{cm}_{c}\right)=H\left(\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}\mathbf{z}_{i}\right).$ (8)
By the convexity of entropy, we have:
$H\left(\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}\mathbf{z}_{i}\right)\leq\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}H(\mathbf{z}_{i}).$ (9)
This inequality implies that the entropy of the centroid prototype is generally lower due to the averaging effect, which reduces the divergence among the samples, leading to lower entropy. Given that $\mathbf{hm}_{c}$ is the sample with the maximum individual entropy among the set $\{\mathbf{z}_{1},\mathbf{z}_{2},\ldots,\mathbf{z}_{N_{c}}\}$, it follows that:
$H(\mathbf{hm}_{c})\geq\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}H(\mathbf{z}_{i})\geq H(\mathbf{cm}_{c}).$ (10)
@@
$\mathcal{L}_{HPCL}^{v}=\frac{1}{N_{v}}\sum_{i\in N_{v}}\frac{-1}{|V_{i}|}\sum_{v\in V_{i}}\log{\frac{\text{exp}\left({\phi}_{0}(v)\cdot{hm}_{+}^{v}/\tau\right)}{\sum\limits_{j\in N_{v}}\text{exp}\left({{\phi}_{0}(v)}\cdot{hm}^{v}_{j}/\tau\right)}},$ (11)
$\mathcal{L}_{HPCL}^{r}=\frac{1}{N_{r}}\sum_{i\in N_{r}}\frac{-1}{|R_{i}|}\sum_{r\in R_{i}}\log{\frac{\text{exp}\left({\phi}_{0}(r)\cdot{hm}_{+}^{r}/\tau\right)}{\sum\limits_{j\in N_{r}}\text{exp}\left({{\phi}_{0}(r)}\cdot{hm}^{r}_{j}/\tau\right)}},$ (12)
$\mathcal{L}_{HPCL}=\mathcal{L}_{HPCL}^{v}+\mathcal{L}_{HPCL}^{r},$ (13)
@@
$hm^{v}_{i,t}=\alpha{hm}^{v}_{i,t-1}+(1-\alpha)\phi_{0}(v),\forall v\in V_{i}$ (14)
$hm^{r}_{i,t}=\alpha{hm}^{r}_{i,t-1}+(1-\alpha)\phi_{0}(r),\forall r\in R_{i}$ (15)